OpenStudy (babynini):

Solve the given equation 2cos^2(theta)+sin(theta)=1

OpenStudy (babynini):

I've gotten up to 2sin^2(theta)+sin(theta)+1=0 but cann't figure out how to factor :/

OpenStudy (babynini):

(2sin(theta) ? ___)(sin(theta) ? ____)

OpenStudy (freckles):

testing what you said for my ownself: \[2(1-\sin^2(\theta))+\sin(\theta)-1=0 \\ \text{ by pythagorean identitiy; also subtract 1 on both sides } \\ 2-2\sin^2(\theta)+\sin(\theta)-1=0 \\ \text{ by distribute property } \\ -2\sin^2(\theta)+\sin(\theta)+1=0 \text{ combining like terms }\] Think you left a sign off on the first term here. I'm also going to multiply -1 on both sides: \[2 \sin^2(\theta)-\sin(\theta)-1=0\]

OpenStudy (babynini):

ah yeah I didn't notice the 2sin^2theta was negative.

OpenStudy (freckles):

anyways if you are doing trial factors you don't have a lot of trials here since -1 is 1(-1) or -1(1)

OpenStudy (babynini):

wait so now what do i do for factoring =.=

OpenStudy (freckles):

if you don't like factoring much you could use the quadratic formula

OpenStudy (babynini):

(2sin(theta) - 1)(sin(theta)+1) = 0 ?

OpenStudy (babynini):

I think factoring is usually nicer xD

OpenStudy (babynini):

dangit. that wouldn't give us the right answer.

OpenStudy (freckles):

well if you want to do factoring if you multiply that out the middle term will be 2sin(theta)-sin(theta which will not be -sin(theta)

OpenStudy (freckles):

so which where the -1 and 1 are

OpenStudy (babynini):

right.

OpenStudy (freckles):

switch *

OpenStudy (babynini):

huh? wouldn't that give the same answer?

OpenStudy (freckles):

\[2 \sin^2(\theta)-\sin(\theta)-1=0 \\ (2 \sin(\theta)+1)(\sin(\theta)-1)=0\]

OpenStudy (babynini):

which would be 2sin^2(theta)-2sin(theta)+sin(theta)-1=0 which is what we want yeah?

OpenStudy (freckles):

yep

OpenStudy (babynini):

I was confused about -2sin(theta) - sin(theta) = -sin(theta) but now I see hah

OpenStudy (freckles):

no no -2sin(theta)+sin(theta)=-sin(theta)

OpenStudy (freckles):

oh did you mean to write that

OpenStudy (babynini):

Shiz yeah, sorry.

OpenStudy (babynini):

Finals week man. Killing me.

OpenStudy (babynini):

K so now we have sin(theta)=1/2 and sin(theta)=1

OpenStudy (freckles):

almost

OpenStudy (babynini):

-1/2?

OpenStudy (freckles):

\[2 \sin^2(\theta)-\sin(\theta)-1=0 \\ (2 \sin(\theta)+1)(\sin(\theta)-1)=0 \\ \sin(\theta)=\frac{-1}{2} \text{ or } \sin(\theta)=1 \] last equation right just a sign off on the first

OpenStudy (babynini):

K and sin = -1/2 at 5pi/6 and 7pi/6 ....right

OpenStudy (freckles):

can I ask if you want me to check your answers can you give me the intervals in which you want to solve them

OpenStudy (babynini):

hm? it just says solve the given equation. No interval.

OpenStudy (freckles):

probably looking for all solutions then

OpenStudy (babynini):

so perhaps 5pi/6, 7pi/6, pi/2

OpenStudy (freckles):

ok but sin(5pi/6)=1/2 not -1/2

OpenStudy (babynini):

crap. 11pi/6

OpenStudy (freckles):

sin(7pi/6)=-1/2 so 7pi/6 is a solution sin(11pi/6)=-1/2 so 11pi/6 is a solution yep yep

OpenStudy (freckles):

and yes sin(pi/2)=1 so pi/2 is a solution to the equation sin(u)=1

OpenStudy (freckles):

but

OpenStudy (freckles):

if they want all the solutions and since we are working with sin and cos just +2pi*n and say where n is an integer

OpenStudy (freckles):

\[\theta=\frac{7\pi}{6}+2 \pi n \\ \theta=\frac{11\pi}{6}+2 \pi n \\ \theta=\frac{\pi}{2}+2 \pi n \\ \text{ where } n \text{ is integer }\]

OpenStudy (babynini):

Ah ok. I remember this :)